OpenStudy (cheesecakekitten):

The sum of two numbers is 14. The larger number minus three times the lesser number is -2. What is the lesser number? A. 10 B. 12 C. 3 D. 4

OpenStudy (zpupster):

you need 2 equations L= larger s= smaller we know L+s=14 and L-3s= -2 solving for L in first L=14-s substitute into 2nd eq. 14-s-3s =-2 solve for s